FEL Upgrade Project Weekly Brief
April 30-May 4, 2001

This was a good week for making progress on many of the hardware systems for the IR Upgrade.  See the WBS reports for the (numerous) details.

The project monthly report for the month of March was distributed earlier this week.

We thank Eli Zimet for releasing an additional $250k of ONR FY01 funding for the Phase 2 upgrade effort.  This brings the project budget to within $50k of the original budget ($4.5M) of the phase 2 work statement.  This document will be adjusted accordingly.

The project team continues to work with the JTO office to firm up the contract and start date for the JTO laser materials studies which we hope begins this month.  We are also working with JTO and JTO’s contractor on the technical planning for the high power FEL workshop planned for June 5-6.

WBS 4 (Injector):
Completed drawings for the shield assembly (door swing mechanism) and submitted to the shop for manufacturing. The ceramic stack pieces have been fitted and stacked, awaiting welding.  Design continued on the interface between the gun and light box, defining an internally tapered tube at the exit of the gun.

WBS 5 (SRF):
Vendor sight visit to Aboloty Engineering to check on the vacuum vessel and space frame fabrications.  Space frame was found out of tolerance and vendor is making plans for straightening vacuum vessel top, but flanges were found out of tolerance and vendor is making plans for a fix.

Vendor site visit to ADC, thermal shield vendor.  Inspected parts, material, QA program, and project planning and integration with other contracts.  All fabricated parts, materials and programs are in good order.

WBS 6 (RF):
Zone 3 - The circulators were tested this past week.

Zone 4 - RF Control Modules were installed temporarily for an experiment to test the stability of the klystrons by adjusting the filament voltage to control klystron beam current.  The missing circulator for position 8 will be installed as time is available.

Injector HVPS - PSC has been given a deadline of next Monday to either provide us a firm price to upgrade their HVPS or no bid our RFP.  I am talking with other companies to provide the various components if the task is moved in-house.

Injector Klystrons - In performing a helium leak test on the collector of the first 100kW klystron, CPI, feels they may have received parts that were not made of the specified oxygen-free copper.  They are making additional tooling to make a definitive test.  This second test is delaying the project 3 weeks.  If they have to remake the parts with new copper, the project will slip 3 to 4 months according to their first estimates.

WBS 8 (I&C):
The drive laser pulse truncator specification is complete and the design is proceeding. Many thanks to Curt Hovater, Trent Allison and the EID group for their design assistance. This device will insert between the existing Drive Laser Pulse Controller (DLPC) and allow for electron beam macropulses (in 500 nanosecond steps) between the existing 1 microsecond and 10 microsecond selections that currently exist. This will also allow for macropulse frequencies to extend beyond the existing 60 Hz to 100 Hz, 300 Hz, 1 kHz, 3 kHz, 10 kHz, and to 20 kHz. During the original FEL design it was doubted that the FEL could even lase in pulsed mode, hence the gap between 1 and 10 microseconds (go figure!). This will be inserted between the DLPC and the ConOptics 305 mode locker controller. There will be not compromise to the MPS since it can only "pull down" the E/O cells and not supply power to gate them ON. This truncator will operate in pulsed, CW and in Alignment mode. The important feature is that it will be synchronous with the FEL output pulses; that is phase stable for experiments like Bob Austin's who require the 18MHz signal to gate the cryodetectors for enhanced noise immunity. As always Al Grippo will create a simple EPICS screen so all can enjoy.
The viewports have been ordered for the upgrade. An order for 40 pcs.; 4.62 flange fused quartz viewports with a clear aperture of 2.69" was placed.  Once these arrive they will have 4 "blind tapped" holes added for mounting the beam viewer camera assembly, then have a gold coating sputtered on. The shop will most likely do the holes and Larry Phillips will supply the Au. A Golay Cell was also ordered, this will function as a spare for the Happek interferometers and also be used to evaluate THz radiation from a port at the 2F region.
A diamond window from ALS was installed at 2F03; the exit of the upstream optical chicane. This port will be used to evaluate the THz radiation emitted from the chicane bend.  The X-Ray detector's, downstream of the wiggler, Pb shield was installed and the control/software was updated (moved over from the scrapers) - the "X" (up/down) & "Y" (right/left) have been adapted to suit the newly reincarnated operational modes required from the backscattered X-Ray experiments.  There was an ion pump VIP2F00A which had been "pitching an occasional fit" that was also changed out.  This ion pump acted in the same fashion as did one on the upstream optical table six moths ago.  The symptoms were occasional trips (to the 5E-7 valve closing interlock threshold). The DSP56F805 microcontroller software is proceeding well.  This device is slated to be used for a number of applications; BPMs, OBPMs, trim PS cards, and for some simple motor control. The Code Warrior software, including compiler and debugging tools have been installed. A sample program has been edited, compiled and loaded with and observed running in the DSP, with variables in memory changing in accordance with the program.  A simple program was also coded to read a built-in A/D channel.  Program appears to work, but have not yet applied a voltage to see the data change.

Progress has been made for remote control of the mirror test stand YAG laser; the initial version of EPICS software for monitor/control is complete and being debugged.  The coding for the (new) VME interface board device driver is complete.  The AVME9510 has all of the control and status functions in a single board (& at a reasonable $); digital in/out and analog in/out.  The field wiring design is being worked out and should be installed next week.  The
LSS connections have been made and tested.  The X-Ray DC Motor has been wired up and is operational through EPICS.  The software control is also completed and being tested for 2 DC motors to be used to control x-ray detector position.
The Lab #2 LSS low level certification passed (although it was only done for training purposes). There exist NO known problems with the LSS master, or any of the labs (except #5 which has not been commissioned).  The new hutch in lab #6 is complete and awaiting a decision on how and where the penetrations should go.  Test boxes were built for testing the Beam Viewer Chassis and the PLC in the LSS system.  The next effort on the LSS is to clean up the overall system print.
There has been more improvements to the website http://laser.jlab.org to include a feedback link, technical document links and more system information.  Each of the systems is getting an overhaul - this is critical to have an accurate "as-built" reflecting all of the changes made over the last couple years prior to the "rip-out".
Cables were installed in the Drive Laser Clean Room to facilitate OBPM testing.  Some problems have been identified with the new boards; the AD549 op-amp was laid-out from top view not bottom view - the damaged chips ere removed and new one soldered in from the bottom of the board. Testing is continuing, this will also act as a platform for system test of the pico motor transport mirror control system.  A turning mirror is installed on a picomotor stage; by offsetting the quad detector the mirror should track.  Twelve boxes have been manufactured in the machine shop and are painted awaiting assembly.

Master machine had the TDR check done on 20 BPM's, these may be complete for delivery next week. Also there are 18 viewers ready to assemble.  We should receive a minimum of 9 viewers next week.

WBS 9 (Transport):
Injector Dipoles (DU/DV)
o  DULY Research is working on adapting their magnetic model of the small injector Dipole (GV) into a
    practical hardware design.
o They delivered the details of the coil for the GU for checking.
Optical Chicane Dipole (DW)
o The design package is almost ready.
Arc Dipoles (GY, GX, GQ)
o We ordered the material for the pole tips of the GY.
o The magnetic analysis of the GY with path length correctors and now synchrotron light periscopes continued.  At
    week’s end David Douglas was given the latest field values along four paths through the magnet.  He was going
    to match his particle pusher analysis with the results found by Tom Schultheiss of AES.
o We continued design and magnetic modeling the Arc Bend Dipole (GX).  We developed the bus bar mounting
    scheme to allow upper and lower halves of the coils to be powered independently and allowed the magnets to
    be flipped over 180° with the same hardware.  Because the magnet is an asymmetrical four sided figure; it is
    used right side up in two locations and flipped over in the remaining two locations.  Our design makes a
    universal magnet that can be used in both orientations.
QG (3 inch quad)
o The coil-manufacturing package remains in Procurement cycle.
o The core drawings received minor modifications in anticipation of signing.  Jeff Karn is starting discussions with
    Procurement on how best of purchase the cores.
QX (Trim Quad)
o We continued debugging the magnetic model going through several iterations with the Vector Fields Co. on how
    to make it yield proper results.
Corrector Dipoles (DB & DJ)
o The package is in vendor’s hands for a re-bid.
o The Dipole Steel continues in Procurement cycle.
o We continued design layout of the girders between cryomodule regions.
o We continued to make advances in the configurations of the arc chambers.

WBS 10 (Wiggler):
Previous magnet measurements showed the field integral in the dispersion section to be non-zero.  A temporary trim coil was wrapped around all poles to provide a "tweakable" offset. Preliminary measurements show such a coil can be used to zero out the integral.  Additional measurements are in the works.  In the wiggler, temporary coils were configured on the field clamps to compensate for the "missing pole."  Magnet measurements showed the extra trim coils could be tweaked to balance pole fields.  G. Biallas is pursuing building production coils.  C. Hutton has resumed the design of the test stand hardware.  It is looking like we should be ready to start wiggler measurements by the fourth week of May.

WBS 11 (Optics):
The design for the ultraviewer is complete; we anticipate the drawing package going out for bid early next week.  Prints for new, near Brewster windows for the User Labs have been signed and are going to Procurement.  The fabricator for the prototype deformable mirrors has proposed several changes that will result in cost savings and accelerated delivery.  (What more can a WBS manager ask for?).  The rest of the internal mounting components for the mirror test stand arrived and are being cleaned.  The manufacturer of our laser interferometer, Veeco, has completed the metrology of the 3" diameter silicon mirrors produced by RMI. They found that the mirrors met or exceeded specifications, qualifying RMI as a source for substrates for the IR Upgrade.  It was also an excellent test of the interferometer, because some of these measurements were near the limits of its capabilities. This gives me confidence it will perform well as part of the Mirror Test Stand.

We set up a pickoff in the Drive Laser Enclosure that allows I & C to test OBPM detectors.  It will also permit testing of the automatic stabilization features of the FEL laser beam before we actually use it.

I had a telephone conference with representatives from NRL regarding the types of instrumentation they are offering to bring to JLab for the Joint Technology Office (JTO) contract

No operations this week.